OpenStudy (anonymous):

Could someone please help me get these two inequalities in slope intercept form so I am able to graph them? Thanks!

OpenStudy (anonymous):

\[x+y \ge 5\] and 2x-y<5

OpenStudy (anonymous):

inequaliy in slope -intercept form is y (inequality sign) m*x +c, where m is the slope and c is the intercept. apply same principle here. can you do the rest?

OpenStudy (shamim):

can u draw a graph by the equation x+y=5

OpenStudy (anonymous):

@Arnab09 I know what slope intercept form is, but I do not know how to put the two inequalities in that form.

OpenStudy (anonymous):

okay, draw the graphs of equations x+y=5 and 2x-y=5 using slope-intercept form. you can do that right?

OpenStudy (shamim):

y=-x+5?

OpenStudy (anonymous):

@shamim I think so. Is that equation considered slope intercept form? If so, wouldn't I just put a 1 in front of the x for my slope? I am probably making this way harder than it is, but I am very confused!

OpenStudy (shamim):

slope intetcept form is y=-x+5 is not it

OpenStudy (shamim):

we wanna draw a graph for this line of equation

OpenStudy (shamim):

can u tell me y=? when x=0